OpenStudy (anonymous):

How much 3M HCl should be used to prepare a 1L solution of pH 1.50?

OpenStudy (aaronq):

use: \(pH=-log[H^+]\) and \(Molarity=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

10^(-1.50) = 0.03162 M H{+} (1 L) x (0.03162 mol/L H{+}) x (1 mol HCl / 1 mol H{+}) / (3.0 mol/L HCl) = 0.01 L